OpenStudy (anonymous):

You want to deposit $12,000 in a bank at an interest rate of 8 percent per year. What is the future value of this money after five years?

Directrix (directrix):

Is the interest compounded continuously or some other way?

OpenStudy (anonymous):

what

Directrix (directrix):

What formula are you supposed to use to compute the future value of the money?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Call a i deposit on a monthly x is the interest rate n is the number of months post we have After January, the amount of a + ax = a (x + 1) fra In early February: a×(x+1)+a=a×(x+1+1)=(ax)[(1+x)2−1] After February: (ax)[(1+x)2−1]+(ax)[(1+x)2−1]x=(ax)[(1+x)2−1]x=(ax)[(1+x)2−1]((x+1)) ...... After n months, the amount of principal and interest are: (ax)(x+1)[(1+x)n−1] Applying the formula, instead of on and